### v1

# Opara: Exploiting Operator Parallelism for Expediting DNN Inference on GPUs

Abstract
GPUs have become the defacto hardware devices to accelerate Deep Neural Network (DNN) inference in deep learning (DL) frameworks. However, the conventional sequential execution mode of DNN operators in mainstream DL frameworks cannot fully utilize GPU resources, due to the increasing complexity of DNN model structures and the progressively smaller computational sizes of DNN operators. Moreover, the inadequate operator launch order in parallelized execution scenarios can lead to GPU resource wastage and unexpected performance interference among operators. To address such performance issues above, we propose Opara, a resource- and interference-aware DNN Operator parallel scheduling framework to accelerate the execution of DNN inference on GPUs. Specifically, Opara first employs CUDA Streams and CUDA Graph to automatically parallelize the execution of multiple DNN operators. It further leverages the resource demands of DNN operators to judiciously adjust the operator launch order on GPUs by overlapping the execution of compute-intensive and memory-intensive operators, so as to expedite DNN inference. We implement and open source a prototype of Opara based on PyTorch in a non-intrusive manner. Extensive prototype experiments with representative DNN and Transformer-based models demonstrate that Opara outperforms the default sequential CUDA Graph in PyTorch and the state-of-the-art DNN operator parallelism systems by up to 1.68×1.68$\times$1.68 × and 1.29×1.29$\times$1.29 ×, respectively, yet with acceptable runtime overhead.
Index Terms:
DNN inference, DNN operator parallelism, scheduling, GPU resource utilization
1 Introduction
Deep Neural Networks (DNNs) have gained notable success in various business fields such as image processing, speech recognition, and virtual reality [1]. In general, DNN inference tasks are exceptionally latency-sensitive. For instance, latency requirements in autonomous driving scenarios are non-negotiable (e.g., within $100$ milliseconds) due to safety considerations [2]. Accordingly, increasing attention from both academia and industry has been paid to efficient model serving. To meet such performance requirements, modern cloud datacenters are hosting thousands of GPUs to accelerate DNN inference for users. For instance, Alibaba Cloud houses more than $6,000$ GPUs, many of which are tasked with managing a substantial volume of inference requests [3].
Cloud-based GPUs are equipped with an increasing amount of computational power, which typically exceeds the resource demands of individual inference tasks, leading to under-utilization and wastage of hardware resources [4]. To achieve the objective model accuracy with fewer computations, several recent works (e.g., Szegedy et al. [5]) focus on substituting large operators with several smaller and multiple-branch operators in DNN models, which further exacerbates the under-utilization of GPU resources. While batching requests [6] or concurrent processing of multiple model inference tasks [7] can mitigate such GPU under-utilization, it inevitably causes increased latency for model inference due to batching latency and performance interference [8]. As a result, it is compelling to improve the GPU utilization without impacting the DNN inference latency. As DNN models can typically be represented by a Directed Acyclic Graph (DAG) with parallel operators, it provides us an opportunity to exploit operator parallelism for accelerating DNN inference on GPUs while improving the GPU utilization.
Unfortunately, it is nontrivial to efficiently parallelize the execution of DNN operators for a DNN inference task due to the following two facts. First, the model DAG typically exhibits considerable complexity, often incorporating hundreds of operators with complex inter-operator dependencies. For simplicity, existing deep learning (DL) frameworks execute DNN operators one by one in topological sorting order [9], which overlooks the parallelization opportunities among operators. To achieve operator parallelism, a recent work (i.e., Nimble [10]) relies on a reduction transformation of the DNN computation graph, which inevitably brings heavy computation overhead. Second, inadequate operator parallel scheduling can adversely impact the DNN inference performance. As evidenced by motivation experiments in Sec. 2.3, the inadequate operator launch order in mainstream DL frameworks (e.g., PyTorch) can prolong the DNN inference latency by up to $29\%$, due to the GPU blocking caused by the non-preemption feature of CUDA kernels [11] and performance interference among parallelizble operators [8]. In addition, several existing works (e.g., IOS [12]) fail to consider the operator launch and function call overhead due to excessive CPU-GPU interactions when parallelizing DNN operators in the DL framework.
To address the challenges above, in this paper, we design Opara, a resource- and interference-aware DNN Operator parallel scheduling framework, with the aim of expediting the execution of DNN inference while improving the GPU utilization. We make the following contributions as below.
[…]
2 Background and Motivation
In this section, we first introduce how DNN operators are executed in mainstream DL frameworks, and identify the key factors that cause the low GPU utilization when serving DNN inference on GPUs. We then conduct motivation experiments to show how to judiciously parallelize the operator executions on GPUs.
2.1 DNN Operator Executions on NVIDIA GPUs
After being scheduled on GPUs, a DNN operator is actually recognized as a kernel. In general, a kernel comprises multiple thread blocks, which are the smallest scheduling granularity in CUDA. A thread block is scheduled to a Streaming Multiprocessor (SM) once the SM has sufficient resources to meet its resource demands [13]. In particular, an SM can concurrently execute multiple thread blocks, and each SM is constrained by a limited number of threads, shared memory, and registers.
To enable parallel executions of operators, we launch operators on multiple CUDA Streams. Each CUDA Stream is actually a task queue that executes tasks sequentially. The execution order of kernels in different CUDA Streams is determined by their arrival order at the stream head. In general, the kernel execution time is considerably short as the batch size is typically small (i.e., ranging from $1$ to $16$) in DNN inference scenarios. Accordingly, the kernel launch overhead constitutes the primary time cost for DNN inference, which offsets the performance gains achieved by the parallel executions of kernels in multiple CUDA Streams. To reduce such overhead, CUDA Graph is a key feature introduced from CUDA 10 that allows scheduling multiple DNN operators on a GPU device at a time.
2.2 Low GPU Utilization Due to Sequential Execution of DNN Operators
Mainstream DL frameworks generally do not support inter-operator parallel executions for DNN inference, due to the complexity of parallel programming. Instead, they execute DNN operators sequentially in topological sorting order, which cannot fully utilize the GPU resource in general. To illustrate that, we conduct motivation experiments using the stock PyTorch 2.0 with GoogLeNet [14], Inception-v3 [5], and BERT [15] on an NVIDIA A100-PCIE-40GB GPU and an NVIDIA RTX 2080 SUPER GPU. In particular, we adopt the SM efficiency [^1] The SM efficiency denotes the proportion of time during which at least one warp is active and executing instructions on a specific SM in a GPU. measured using NVIDIA Nsight Compute CLI [^2] Nsight Compute CLI: https://docs.nvidia.com/nsight-compute/NsightComputeCli/index.html to evaluate the GPU utilization.
Figure 1: Average SM efficiency of an A100-PCIE-40GB GPU when running GoogLeNet, Inception-v3, and BERT models with different batch sizes.
Figure 1: Average SM efficiency of an A100-PCIE-40GB GPU when running GoogLeNet, Inception-v3, and BERT models with different batch sizes.
Figure 2: Inference latency of GoogLeNet running on an RTX 2080 Super GPU with different operator launch orders and batch sizes.
As shown in Fig. 2, DNN inference on the mainstream DL framework (i.e., the latest PyTorch 2.0) leads to relatively low GPU utilization on A100. Specifically, the SM efficiency of GoogLeNet with the batch size set as $1$ is a mere $2.53\%$, while that of Inception-v3 is $12.04\%$. Even when the batch size increases to $16$, the SM efficiency of GoogLeNet is still $9.12\%$ and that of Inception-v3 reaches $21.48\%$. Similarly, larger models such as BERT cannot fully utilize the GPU resource. By setting the sequence length as $32$, the SM efficiency of BERT with the batch size set as $1$ and $16$ is $18.5\%$ and $54.8\%$, respectively. Meanwhile, we repeat our experiments on a less powerful GPU (i.e., RTX 2080 SUPER), and the SM efficiency of the three workloads ranges from $10.47\%$ to $82.98\%$. Our experiment results indicate that the sequential execution of DNN operators is the root cause of low GPU utilization for running DNN inference, and exploiting the operator parallelism can expedite the DNN inference on GPUs while improving GPU utilization.
2.3 Performance Impacts of Operator Launch Order
Apart from the sequential execution of DNN operators, the inadequate operator launch order (i.e., the topological sorting order of the model DAG) in mainstream DL frameworks can also lead to idle GPU resource usage and performance interference, thereby prolonging the inference latency. We conduct two motivation experiments to identify why the operator launch order can impact the inference latency.
GPU Blocking. As each operator has a different number of blocks requiring three types of resources for execution, i.e., threads, shared memory, and registers, a resource-unaware operator launch order can easily block the execution of operators until enough resources become available on the GPU. Such GPU blocking can severely waste the available GPU resources. As shown in Fig. 2, changing the operator launch order from order $1$ (i.e., depth-first topological sorting) to order $2$ (i.e., Opara designed in Sec. 3) for GoogLeNet can reduce the inference latency by up to $29\%$ with different batch sizes. Furthermore, we repeat such an experiment on the A100 GPU, and the experiment results show around $10.3\%$ of performance improvement by optimizing the operator launch order for GoogLeNet.
Figure 3: Overlapping the execution of compute-intensive and memory-intensive operators denoted as red circles and blue circles, respectively.
Performance Interference. As the performance interference among operators can prolong the inference latency [8], we further conduct another experiment on A100 to illustrate the effectiveness of overlapping the execution of compute-intensive and memory-intensive operators in mitigating the inference. As depicted in Fig. 3 (case 1), prioritizing the parallel execution of Relu and Conv operators can cause less severe interference, compared with parallelizing two Relu operators, leading to a $13.6\%$ reduction in the inference latency. Similarly, prioritizing the launch order of Add operator in case 2 can increase the inference performance by $12.7\%$, simply because the execution of compute-intensive and memory-intensive operators is overlapped.
Summary. Low GPU utilization of DNN inference is mainly caused by two factors: First, the sequential execution of DNN operators cannot fully utilize the GPU resources. Second, the default topological sorting order of operator launch is commonly resource- and interference-unaware. Accordingly, judiciously parallelizing the DNN operators with an adequate operator launch order can accelerate DNN inference on GPUs while improving the GPU utilization.
3 System Design
In this section, we design Opara illustrated in Fig. 4, an operator parallel scheduling framework to reduce DNN inference latency while improving the GPU resource utilization. Specifically, Opara takes DNN models and input tensors (i.e., inference data) from users. According to the operator dependencies in the model DAG, the Stream Allocator first employs a stream allocation algorithm to determine which stream the operators should be allocated to. The Model Profiler then gathers the resource demands of each operator using the model profiling. With such resource demands of operators, the Operator Launcher further employs a resource- and interference-aware operator launch algorithm to optimize the operator launch order on GPUs. Finally, the Graph Capturer generates a parallelized CUDA Graph by combing the stream allocation plan and operator launch order, thereby enabling efficient DNN inference on GPUs.
Figure 4: System overview of Opara.
3.1 Stream Allocator
To parallelize the execution of operators in CUDA Streams, we leverage the computation graph (i.e., DAG) of DNN models to determine how many streams to launch and how to allocate operators to the streams.
Definition of A Model DAG. DNN computation graph can be represented as a DAG $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of operators in the model, and $\mathcal{E}$ denotes the operator dependencies. Each vertex ${v}\in\mathcal{V}$ denotes a DNN operator (e.g., Conv, MaxPool). Each edge $\langle u,v\rangle\in\mathcal{E}$ denotes the operator dependency, where $u$ is a predecessor of ${v}$ and ${v}$ is a successor of ${u}$. The set of all predecessors of an operator ${v}$ are denoted as $\mathcal{N}_{pred}$. The set of all successors of an operator ${v}$ are denoted as $\mathcal{N}_{succ}$.
Problem Formulation and Analysis. As a maximum of $|\mathcal{V}|$ streams can be launched for a DAG, we simply use a matrix $\mathcal{A}$ of size $|\mathcal{V}|\times|\mathcal{V}|$ to represent the stream allocation plan. Each element $a_{ij}\in\mathcal{A}$ is a boolean value, indicating whether the $i$-th operator is executed in the $j$-th stream. Accordingly, a stream allocation plan $\mathcal{A}$ can determine how the DNN operators are parallelized and synchronized. The inference latency $T_{inf}$ of a DNN model is formulated as
$T_{inf}=T_{para}+T_{overhead},$ (1)
where $T_{para}$ denotes the parallelized execution time of a DNN model and $T_{overhead}$ denotes the operator synchronization overhead given a stream allocation plan. In more detail, as the stream allocation plan $\mathcal{A}$ can parallelize the execution of DNN operators, $T_{para}$ can be formulated as
$T_{para}=h(\mathcal{A})\times T_{seq},$ (2)
where $T_{seq}$ denotes the inference latency of sequential execution of DNN operators and $h(\mathcal{A})\in(0,1]$ denotes the inference acceleration factor achieved by operator parallelism with the stream allocation plan $\mathcal{A}$. In particular, a certain number of synchronization operators need to be inserted into the model DAG, to ensure the parallelized execution of the model. Such a process introduces the operator synchronization overhead $T_{overhead}$, which is formulated as
$T_{overhead}=g(\mathcal{A})\times t_{overhead},$ (3)
where $t_{overhead}$ is the time overhead caused by one synchronization operator and $g(\mathcal{A})$ is positively correlated with the number of synchronization operators in approximation.
By substituting Eq. (3) and Eq. (2) into Eq. (1), we aim to minimize the inference latency $T_{inf}$ by identifying an optimal stream allocation plan $\mathcal{A}$. Accordingly, we formulate the stream allocation optimization problem as
$\min_{\mathcal{A}}$ $T_{inf}=h(\mathcal{A})\times T_{seq}+g(\mathcal{A})\times t_{overhead}$ (4)
s.t. $\sum_{j=1}^{|\mathcal{V}|}a_{ij}=1,\qquad\forall i\leq|\mathcal{V}|$ (5)
[…]
0:  DNN computation graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
0:  Set of streams to be launched $\mathcal{S}$ used by the operator set $\mathcal{V}$.
1:  Initialize: $\mathcal{S}\leftarrow\emptyset$;
2:  for each operator $v\in\mathcal{V}$ do
3:     for each predecessor $p\in\mathcal{N}_{pred}$ of $v$ do
4:        if $v$ is the first successor of $p$ then
5:           stream of $v\leftarrow$ stream of $p$;  // put $v$ and $p$ in the same stream
6:           break out of the loop;
7:        end if
8:     end for
9:     if $stream$ of $v$ is $null$ then
10:        $stream$ of $v\leftarrow$ launching a stream;  // put $v$ in a newly launched stream
11:        $\mathcal{S}\leftarrow\mathcal{S}\cup\left\{stream\text{ of}\>v\right\}$;
12:     end if
13:  end for
14:  return  $\mathcal{S}$.
Algorithm 1 Stream allocation algorithm in Opara.
Stream Allocation Algorithm. To solve such a complex stream allocation problem in polynomial time, we design a heuristic algorithm in Alg. 1. The key idea of our algorithm is to allocate parallelizable operators to multiple CUDA Streams as much as possible (i.e., minimizing the value of $h(\mathcal{A})$). To avoid excessive synchronization operations (i.e., minimizing the value of $g(\mathcal{A})$), we aim to greedily put non-root nodes (i.e., operators) in the same CUDA stream as one of their predecessor operators. Specifically, given a computation graph $\mathcal{G}$, Opara first initializes a set of streams to be launched $\mathcal{S}$ and then enumerates operators in $\mathcal{V}$ in topological sorting order (lines $1$-$2$). For each operator ${v}\in\mathcal{V}$, it iterates over all of its predecessors $p\in\mathcal{N}_{pred}$ (line $3$). If $v$ is the first successor of the current predecessor $p$, it allocates $v$ to the same stream of $p$; otherwise, it moves on to the next predecessor (lines $4$-$7$). If $v$ does not find a predecessor that satisfies such a condition above, we allocate the operator $v$ to a newly launched stream (lines $9$-$11$). In particular, the streams consume GPU resources only when operators are running, which indicates that performance interference among streams does not occur when operators are not executed on GPUs.
3.2 Model Profiler
As discussed in Sec. 2.3, the blocks in an operator execute the same instructions even with different data, which indicates that the GPU resources required by the blocks in an operator are the same. Accordingly, we obtain the resource demands of each operator by simply profiling the resource consumption (i.e., the amount of shared memory, the number of registers and GPU threads) of a block in an operator. Such resource demands of operators will be used by Operator Launcher to determine an adequate operator launch order. In particular, we implement our Model Profiler utilizing the torch.profiler.profile() API, and it requires profiling each DNN inference only once to acquire the resource demands information for each operator, thereby bringing acceptable profiling overhead. We will examine the inference profiling overhead of Opara in Sec. 5.3.
3.3 Operator Launcher
Problem Analysis. As illustrated in Sec. 2.3, inadequate operator launch orders can significantly affect the DNN inference latency. To identify an optimal launch order, a naive solution is iterating through all possible topological sorting orders of a model DAG and choose the order with the lowest inference latency. However, such a method involves selecting nodes with zero indegree and deleting the corresponding vertices and their connected edges. By assuming that $n$ operators exist in a model DAG, the time complexity of traversing all topological sorting orders is $\mathcal{O}(n!)$, which is also an NP-hard problem. As a result, we turn to designing a heuristic operator launch algorithm to solve such a complex problem.
0:  DNN computation graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
0:  Operator queue $\mathcal{Q}$ in resource- and interference-aware operator launch order.
1:  Initialize: List of operators to be launched $\mathcal{L}\leftarrow\emptyset$, $\mathcal{L}_{mem}\leftarrow\emptyset$, $\mathcal{L}_{comp}\leftarrow\emptyset$, and $\mathcal{Q}\leftarrow\emptyset$;
2:  Add the operators $v\in\mathcal{V}$ with an indegree of 00 that is memory-intensive and compute-intensive to $\mathcal{L}_{mem}$ and $\mathcal{L}_{comp}$, respectively;
3:  while $\mathcal{L}_{mem}$ or $\mathcal{L}_{comp}$ is not empty do
4:     $\mathcal{L}$ = alternately choose a non-empty list from ${\{\mathcal{L}_{mem},\mathcal{L}_{comp}\}}$;
5:     $v_{min}\leftarrow$ the operator that requires the least amount of GPU resources in $\mathcal{L}$;
6:     .$\mathcal{L}.{remove}({v_{min}})$, and $\mathcal{Q}.{append}({v_{min}})$;   // launch the operator $v_{min}$
7:     for each successor $s\in\mathcal{N}_{succ}$ of $v_{min}$ do
8:        $indegree$ of ${s}\leftarrow$ $indegree$ of ${s}-1$;   // update the $indegree$ of ${s}$
9:        if $indegree$ of ${s}==0$ then
10:           if operator ${s}$ is memory-intensive then
11:              $\mathcal{L}_{mem}.{append}({s})$;  // add $s$ to $\mathcal{L}_{mem}$
12:           else
13:              $\mathcal{L}_{comp}.{append}({s})$;  // add $s$ to $\mathcal{L}_{comp}$
14:           end if
15:        end if
16:     end for
17:  end while
18:  return  $\mathcal{Q}$.
Algorithm 2 Operator launch algorithm in Opara.
Resource- and Interference-Aware Operator Launch Algorithm. Launching operators with heavy resource demands first to the GPU is likely to cause resource fragmentation, hindering the GPU executions of subsequent operators. Moreover, the GPU can thus be blocked due to the non-preemptive feature of kernel execution [11]. To mitigate such a problem and maximize the GPU utilization, we design a heuristic algorithm in Alg. 2 to greedily prioritize launching the operators with the least amount of GPU resource demands. Accordingly, it can maximize the parallel executions of multiple operators within a single model. To further mitigate the performance interference among operators [8], we aim to overlap the execution of computation-intensive operators and memory-intensive operators, as classified by our offline-collected operator table. Specifically, it first initializes and maintains a priority queue $\mathcal{Q}$ of operators in resource- and interference-aware operator launch order (line $1$). It then retrieves all operators to be launched with an indegree of 00 in a list $\mathcal{L}$, which alternates between a non-empty list of memory-intensive operators $\mathcal{L}_{mem}$ and a non-empty list of compute-intensive operators $\mathcal{L}_{comp}$ (lines $2$-$4$). Each time the operator requiring the least amount of GPU resources (e.g., shared memory, threads, registers) is chosen from $\mathcal{L}$ and then put into the queue $\mathcal{Q}$ (lines $5$-$6$). In particular, the potential GPU blocking issue faced by remaining large operators is noncritical in our scenario, as the operator list $\mathcal{L}$ is dynamic and can be compensated for the upcoming small operators to be launched. Finally, the operator list $\mathcal{L}_{mem}$ and $\mathcal{L}_{comp}$ are continuously updated by adding new operators with an indegree of 00 (lines $7$-$16$).
3.4 Graph Capturer
To eliminate the overhead caused by kernel launches and function calls, the Graph Capturer first sets the CUDA Streams obtained from the Stream Allocator to the capture mode, and then it launches the operators of the DNN model to these streams according to the operator launch order specified by the Operator Launcher. To ensure the dependencies among operators, the Graph Capturer also launches the necessary synchronization operators to the streams. Consequently, a CUDA Graph is generated to enable operator parallelization while improving the GPU utilization. Such a graph capture process is lightweight and non-intrusive to PyTorch, as it has been exposed as a high-level API in PyTorch officially. We simply use the PyTorch API to capture and then generate the CUDA Graph.
4 Implementation of Opara
We implement a prototype of Opara with around $1,000$ lines of Python codes, which have been integrated into PyTorch 2.0 as a plug-in module. The source codes are currently publicly available on GitHub (https://github.com/icloud-ecnu/Opara). Specifically, we employ torch.fx.Graph as the computation graph for DNN models in Opara. Its Intermediate Representation (IR) allows us to schedule DNN operators directly in the Python environment. In more detail, we leverage the torch.cuda.set_stream() API in PyTorch to launch operators on the CUDA Streams. In particular, as DNN operators are executed asynchronously on multiple streams for parallelized execution, appropriate operator synchronizations are required to guarantee that the parallelized computation conforms to the operator dependencies. We implement the operator synchronizations using the event.record() and stream.wait_event(event) APIs. Finally, we use torch.cuda.graph(g) to generate a CUDA Graph that can execute DNN operators in parallel based on the CUDA Streams. In summary, we build our prototype of Opara only using the high-level APIs of PyTorch in a lightweight manner, rather than modifying the computation graph construction module as in Nimble [10]. Accordingly, Opara is non-intrusive to the DL framework, and it can stably work as long as the framework APIs are not updated.
5 Performance Evaluation
In this section, we carry out prototype experiments to demonstrate the efficacy and runtime overhead of Opara in comparison to the stock PyTorch and state-of-the-art operator parallelism frameworks.
5.1 Experimental Setup
Hardware Configuration and Workloads. We conduct our experiments on an NVIDIA A100-PCIe-40GB GPU with Intel(R) Xeon(R) Gold 6240 CPU @ 2.60GHz, and an NVIDIA GeForce RTX 2080 SUPER-8GB GPU with an Intel(R) Core(TM) i9-10920X CPU @ 3.50GHz. We implement Opara using PyTorch 2.0, CUDA 11.7, and cuDNN 8.5.0, as a plug-in module of PyTorch, as discussed in Sec. 4. Our experiments employ representative DNN models, including Inception-v3 [5], GoogLeNet [14], BERT [15], and T5 [17].
Baselines and Metrics. We compare DNN inference performance achieved by Opara with that achieved by the stock PyTorch (with CUDA Graph disabled), default sequential CUDA Graph, and Nimble [10]. Specifically, Nimble transforms the computation graph into a bipartite graph and then identifies its maximum matching to determine an appropriate stream for each operator. In particular, we focus on four key metrics including DNN inference latency, SM efficiency, and GPU memory consumption, as well as DNN inference throughput. All the experiment results are averaged over $1,000$ runs.
Figure 5: (a) Relative speedup, (b) SM efficiency and peak memory consumption of GPUs running representative DNN models with batch size set as $1$ achieved by PyTorch, CUDA Graph, Nimble, and Opara operator scheduling mechanisms.
5.2 Effectiveness of Opara
End-to-end Inference Latency. We first examine whether Opara can accelerate the DNN inference. As shown in Fig. 5, Opara consistently outperforms the other three baselines with four representative DNN models [^3] As BERT and T5 are GPU memory-intensive, we execute the two models on the A100, while the other two models are executed on the RTX 2080 SUPER.. Specifically, Opara can achieve 1.85×1.85$\times$1.85 × to 4.18×4.18$\times$4.18 × speedup compared to the stock PyTorch. This is because Opara utilizes CUDA Graph to eliminate the operator launch and function call overhead. Opara surpasses the default CUDA Graph by up to 1.68×1.68$\times$1.68 ×, simply because of the parallel execution of DNN operators in Opara. Furthermore, Opara outperforms Nimble by up to 1.29×1.29$\times$1.29 × because it judiciously alternates the scheduling of different types of operators with the lowest GPU resource consumption for each kernel launch time. Moreover, Opara initiates enough streams to increase parallelism (e.g., $28$ streams with Opara versus $4$ streams with Nimble for GoogLeNet), thereby maximizing the operator parallelism.
GPU Utilization and Memory. To unveil the performance gains of Opara, we proceed to look into the GPU utilization (i.e., SM efficiency of GPUs) and memory consumption during the execution of DNN models. As shown in Fig. 5, Opara exhibits a similar improvement in GPU utilization compared to the three baselines as in Fig. 5. Specifically, Opara significantly improves the GPU utilization compared to the stock PyTorch, because Opara mitigates the scheduling overhead of the stock PyTorch. When compared with the default CUDA Graph, Opara increases the GPU utilization of Inception-v3, GoogLeNet, BERT, and T5 by $36\%$, $58\%$, $20\%$, and $19\%$, respectively. Such performance gains come from the parallelized execution of operators. When compared to Nimble, Opara boosts the GPU utilization by 1.01×1.01$\times$1.01 × to 1.42×1.42$\times$1.42 × mainly due to the two facts: (1) maximizing stream allocations in Opara can increase operator parallelism opportunities, and (2) optimizing the operator launch order in Opara further minimizes the GPU idle time. Furthermore, the parallel execution of operators requires an increased amount of data to reside in GPU memory simultaneously, thereby leading to a higher peak GPU memory consumption of Opara than that of sequential executions (i.e., PyTorch and the default CUDA Graph).
Figure 6: Timeline of operator executions during a segment of inference process of GoogLeNet achieved by CUDA Graph, Nimble, and Opara.
Timeline of Operator Executions. To understand the efficacy of Opara, we further illustrate the operator execution timeline by taking a segment of inference process of GoogLeNet as an example. In particular, we leverage NVIDIA Nsight System CLI [^4] Nsight System CLI: https://docs.nvidia.com/nsight-systems/UserGuide/index.html to track the operator executions. As depicted in Fig. 6, we observe that there are $4$ parallelizable operators numbered from $1-4$ during the segment of inference process. CUDA Graph executes these operators sequentially in a stream, and accordingly only operators $4$ and $1$ are executed during the time period. The other two operators (operators $2$ and $3$) are forced to queue up, which leads to a long execution time. Though Nimble can parallelize operators in the order of $1$, $2$, $3$, and $4$, it only schedules $2$ operators on two streams, causing a long GPU idle time. In contrast, Opara prioritizes operator $4$ and initiates more streams than Nimble, so that operators $4$, $1$, and $2$ can be executed at the same time and the operator parallelism is maximized. Accordingly, Opara can achieve the shortest inference latency by exploring operator parallelism compared with CUDA Graph and Nimble.
(a) T5
(b) BERT
Figure 7: Timeline of Transformer-based models achieved by Nimble and Opara operator scheduling mechanisms. Red rectangles and blue rectangles represent compute-intensive operators and memory-intensive operators, respectively.
Effectiveness of Opara on Transformer-Based Models. We conduct experiments with T5 and BERT model, and Opara outperforms Nimble by $9.3\%$ for the T5 model as shown in Fig. 5. This is because Opara optimizes the launch order of operators in T5 and schedules them into $6$ streams compared with $3$ streams in Nimble. Moreover, the operator diversity in T5 offers Opara overlap the compute-intensive Arange operators and the memory-intensive To and Ones operators, as shown in Fig. 7(a). In contrast, Opara achieves similar operator launch order and the same number of streams as Nimble as depicted in Fig. 7(b). This is because the parallelizable operators of BERT are always the Embedding operators or the Sgemm operators, which reduces the opportunity of operator overlapping and launch order optimization. Accordingly, Opara achieves marginal performance gains for BERT compared with Nimble, yet 1.08×1.08$\times$1.08 × to 4.06×4.06$\times$4.06 × speedup compared to the stock PyTorch and CUDA Graph as shown in Fig. 5.
Figure 8: Inference throughput of Inception-v3 with PyTorch, CUDA Graph, Nimble, and Opara by varying the batch size from $1$ to $32$.
Figure 8: Inference throughput of Inception-v3 with PyTorch, CUDA Graph, Nimble, and Opara by varying the batch size from $1$ to $32$.
Figure 9: Relative speedup of Inception-v3 with PyTorch, CUDA Graph, Nimble, and Opara by varying the batch size from $1$ to $8$.
Throughput under Different Batch Sizes. We further examine the improvement of Opara in terms of DNN inference throughput across varying batch sizes. As depicted in Fig. 9, we observe that Opara consistently surpasses the other three baselines by varying the batch size from $1$ to $32$. Nevertheless, the performance gains of Opara gradually diminish as the batch size increases. As an example, Opara outperforms the default CUDA Graph by 1.41×1.41$\times$1.41 × and 1.09×1.09$\times$1.09 × when the batch size is $1$ and $32$, respectively. This is because the amount of GPU resources occupied by a single operator increases when dealing with larger batch sizes, which results in less amount of GPU resources available for the execution of parallelized operators. The results above also show that maximizing the operator parallelism can expedite DNN inference while improving the GPU utilization.
Effectiveness of Opara on GPUs with Sufficient Resources. To validate the efficacy of Opara on GPUs equipped with powerful computing resources, we repeat experiments by comparing performance gains with the three baselines on the A100. As shown in Fig. 9, we observe that Opara consistently outperforms the three baselines by varying the batch size from $1$ to $8$, mainly due to the better performance of operator parallelism on the A100. In more detail, both Opara and Nimble can benefit from the operator parallelism compared to the stock PyTorch and default CUDA Graph. Moreover, Opara can achieve 1.08×1.08$\times$1.08 × to 1.15×1.15$\times$1.15 × speedup compared to Nimble on the A100, which are slightly smaller performance gains compared to that on the RTX 2080 SUPER. This is because sufficient resources are available on the A100, resulting in GPU under-utilization of DNN inference and limited optimization opportunities for the operator launch order. As the batch size increases to $8$, a higher GPU resource demands of the operator enlarge the optimization space for adjusting the operator launch order, leading to higher performance gains of Opara compared to Nimble.
5.3 Runtime Overhead of Opara
We evaluate the runtime overhead of Opara in terms of algorithm computation time and inference profiling overhead. We first compare the computation time of the stream allocation algorithm in Opara and Nimble [10]. As listed in Table I, Opara incurs algorithm computation time of $0.58$ ms, $0.27$ ms, $0.5$ ms, and $2.8$ ms for BERT, GoogLeNet, Inception-v3, and T5, respectively. In contrast, such computation time in Nimble is $20.8$ ms, $5.80$ ms, $14.40$ ms, and $161.4$ ms for the four models, respectively, which is at least a magnitude larger than that of Opara. This is because Nimble requires a graph transformation together with an exhaustive search in the bipartite graph. Such a process is time-consuming with a time complexity in the order of $\mathcal{O}({n}^{3})$, where $n$ is the number of operators in a model DAG. In contrast, the time complexity of Opara can be reduced to the order of $\mathcal{O}({n})$. This is because the inner loop of the stream allocation algorithm in Opara only depends on the maximum width of the computation graph, which is typically small (i.e., below $20$). As DNN models become increasingly complex and memory-intensive in future [18], the number of operators will also grow exponentially. Such an algorithm overhead in Nimble becomes unacceptable when $n$ is large enough. In addition, as the Model Profiler needs to run the DNN inference only once, Opara requires several (i.e., $4.25$) milliseconds of profiling overhead in our experiment. In sum, the runtime overhead of Opara is practically acceptable.
TABLE I: Computation time (in milliseconds) of the stream allocation algorithm (i.e., Alg. 1) in Opara and Nimble with different DNN models.
BERT GoogLeNet Inception-v3 T5
Opara $0.58$ $0.27$ $0.50$ $2.8$
Nimble [10] $20.8$ $5.80$ $14.40$ $161.4$
6 Related Work
Inter-operator Parallelism within A Single Model. To parallelize the execution of DNN operators, Rammer [19] proposes the concept of “rtask” to allow fine-grained computing scheduling on a GPU device. It enables task combinations in different operators executed on SMs. To maximize the operator parallelism, Cocktailer [20] further proposes to co-scheduling control flow operators and data flow operators into a GPU device based on Rammer. Nimble [10] leverages the bipartite graph algorithm to adequately launch operators on CUDA streams. A recent work (i.e., IOS [12]) deploys operator fusion and dynamic programming to determine operator parallelization plans. However, Nimble and IOS require a lengthy search process and neglect the optimization space of operator launch order as well as the operator launch overhead. In contrast, Opara utilizes the CUDA Graph to eliminate such performance overhead. It also employs a lightweight stream allocation algorithm to achieve inter-operator parallelism. To reduce the GPU idle time and interference, Opara determines a feasible operator launch order according to operator resource demands.
[…]
Intra-operator Parallelism. Existing DL frameworks, such as PyTorch and TensorFlow, employ expert-optimized operator libraries to accelerate individual operators. To reduce human interactions, TVM [24] uses machine learning methods to automatically search for efficient codes. Such a search process is time-consuming and requires artificially-specified parameter space. To achieve fully automated code generation, Ansor [25] implement an automatic search space construction. As the number of compute units in commodity GPUs increases, an individual DNN operator cannot fully utilize the GPU resources. Accordingly, the stream allocation and operator launch algorithms in Opara can work with the intra-operator parallelism methods above to further improve the GPU resource utilization.
7 Conclusion and Future Work
This paper presents the design and implementation of Opara, a lightweight operator scheduling framework to speed up DNN inference on GPUs. By reducing the synchronization overhead among operators, Opara designs a stream allocation algorithm to automatically allocate operators without dependencies to different CUDA streams, thereby achieving operator parallelism effectively. Furthermore, Opara leverages non-intrusive inference profiling to judiciously select an appropriate operator launch order to mitigate interference and maximize the GPU utilization. Extensive prototype experiments show that Opara can improve the performance of DNN inference by up to $29\%$, as compared to the state-of-the-art operator parallelism systems.
We plan to extend Opara in the following directions: (1) constructing an analytical model to analyze the inference latency and the performance interference caused by inter-operator parallelism, and (2) examining the effectiveness of Opara when accelerating more large language models (e.g., GPT-3), as well as (3) implementing Opara on top of other DL frameworks, e.g., TensorFlow.